OpenStudy (anonymous):

helppp! solve simultaneous equation: x=2y-2 x^2=y^2+7?

OpenStudy (anonymous):

you just have to substitute the numbers into x. Because it already said x=2y-2, that means: (2y-2)^2=y^2+7 so you have to do the exponent first: (2y)²=4y² (-2)²=4 So now you will have: 4y^2+4=y^2+7 and now you rearrange the unknowns and the numbers: 4y^2-y^2=7-4 and add/subtract: 3y^2=3 now you want the unknown to be by itself so you have to take of the other numbers with the unknown step by step. So first, take of the 3 by divide it: y^2=1 And then square root it: y=√1 and the answer can be -1 or 1 :D

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

no probs ^^